# ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers

### v2

@@ 5 Conclusions
In this paper, we introduced ESC, the first fully transformer-based neural speech codec designed for multilingual speech coding. ESC surpasses the existing state-of-the-art time-domain codecs in computational efficiency and achieves comparable compression performance even without a powerful adversarial discriminator. Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are more efficient for neural speech coding than the traditional convolutional blocks and residual vector quantization. Overall, our study suggests a promising direction for speech compression. Future research can focus on multi-scale quantization techniques and explore additional transformer variants suitable for speech signal modeling.
@@ 6 Limitations
First, recent audio codecs are increasingly used in downstream generation tasks, where the codec acts as a foundational model to create discrete acoustic representations Borsos et al. (2023); Kreuk et al. (2022); Siuzdak (2023); Wang et al. (2023); Du et al. (2024). These compressed representations are treated as audio tokens and can be used for language modeling, making them more suitable for generation tasks. Our work does not address this significant aspect. Therefore, a promising future direction is to re-evaluate the ESC models in similar downstream applications. We believe that the cross-scale representations learned from transformer backbones would offer advantages over the fixed-scale features of convolutional codecs when applied to speech synthesis tasks. Second, different automatic metrics for audio evaluation often produce inconsistent results, which is also evidenced in our experiments. To strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014). Third, given the scalability of transformers Kaplan et al. (2020), further increasing the ESC model size and training it on more extensive audio datasets is another direction to enhance compression performance. Lastly, as discussed in previous sections, the proposed cross-scale residual vector quantization scheme requires partial decoder layers during compression, which introduces additional latency as the bitrate increases. Future research could focus on refining this scheme and developing alternative multi-scale quantization structures that are more computationally efficient.
@@ Appendix A Pre-training Paradigm
The details of the proposed pre-training paradigm for efficiently learning VQ-based codecs are depicted in Algorithm 3. The training includes a warm start by-passing all the quantization modules and updating encoder and decoder layers only, using the reconstruction losses. In implementation, we set the quantization output $\bm{z}_{q}$ to be equal to the encoded representation $\bm{z}_{e}$. We also zero out the VQ losses to prevent codebook updates.
@@
2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\text{STFT}(\bm{x})))$
3:     $\mathcal{L}=\mathcal{L}_{recon}(\bm{x},\text{I-STFT}(\hat{\mathcal{X}}))$
@@ B.1 DAC Reproduction Setups
Our customized reproduction of DAC models largely follows their released development codes. The original DAC model, designed for 16kHz audio signals, uses 12 VQs, supporting bitrates from 0.5kbps to 6.0kbps. To compare with ESC at the same bitrate levels, we extended the number of codebooks in its RVQ to 18, creating DAC-Base. For DAC-Tiny, we reduced the encoder dimension from 64 to 32 and the decoder dimension from 1536 to 288, while keeping other parameters unchanged. All these DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our dataset. All other configurations can be referred to from the official release [^5] DAC’s official configuration is available here..
@@ B.2 ESC Architecture Configurations
The configurations for ESC-Base are detailed in Table 3. The layer depth indicates the number of STBs cascaded at each encoder/decoder layer. We use GELU activation functions and LayerNorm for normalization. During quantization, our codec processes two overlapped spectrum frames together (i.e., the flattened spectrum feature is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}\rightarrow\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$). Overall, the vector quantizer of ESC-Base at each bitstream uses $\log_{2}1024\times 3\times 150=4500$ bits for every 3-second input speech (1.5kbps). For ESC-Large, we increase the layer depth to 4 while keeping other settings the same. In the variant ESC-Base (Adversarial), the GAN discriminator intervenes in training only after the pre-training stage ends. The GAN discriminator follows the same architecture used in DAC trainings. All three variants are trained in distributed setups for 0.4 million iterations on 4 NVIDIA RTX 4090 GPUs with a batch size of 36, taking approximately 150 GPU hours.
B.3 Details on Ablation Experiments
All ablation models were trained for 0.25 million iterations. The ablation model SwinT+CS-RVQ is exactly our ESC-Base codec. The pre-training stages constitute for 0.025 million iterations. The Swin Transformer encoder and decoder configurations follow those in ESC-Base, as demonstrated in Table 3. CS-RVQ ablation codecs follow the same configuration in Table 3 as well. For CNN-based codecs, their channel dimensions are set to match those of SwinT-based codecs. Each CNN block comprises a residual unit and a downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency dimension. The residual unit consists of two 2D convolutional layers, with a BatchNorm layer and Parametric ReLU activations interleved between them. For codecs based on RVQs, we adapted the vanilla RVQs commonly used in time-domain codecs. To handle 2-dimensional spectrum latents, we perform RVQ within different sub-channels. Specifically, the flattened feature frame is first split into sub-group vectors, similar to product VQs. Each sub-group vector is recursively quantized, as in RVQs. We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5kbps per bitstream, 6 in total). Other training configurations are the same as ESC-Base. All these ablation experiments together took about 100 GPU hours on 4 RTX 4090 GPUs.
@@ B.3 Details on Ablation Experiments
Table 3: Parameter configurations of codec variant ESC-Base, which comprises 6 encoder/decoder layers.

### v3

@@ 5 Conclusions
In this paper, we introduce ESC, the first fully transformer-based neural speech foundation model designed for multilingual speech coding. ESC surpasses existing state-of-the-art time-domain VQ-based codecs in terms of complexity and achieves comparable compression performance without the need for a powerful adversarial discriminator. Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are better suited for neural speech coding than the convolutional blocks and residual vector quantization utilized in mainstream codecs. Overall, our study suggests a promising direction for speech foundation models. Future research could focus on expanding multi-scale vector quantization techniques and investigating additional transformer variants optimized for speech signal modeling.
@@ 6 Limitations
First, recent neural audio codecs are increasingly utilized in downstream generation tasks, where the codec acts as a foundation model to create discrete acoustic representations Borsos et al. (2023); Kreuk et al. (2022); Siuzdak (2023); Wang et al. (2023); Du et al. (2024). These compressed representations, treated as acoustic tokens, are suitable for auto-regressive language modeling in generative tasks. However, our work does not explore this important aspect. A promising future direction would be to evaluate ESC in downstream applications such as speech synthesis and speech recognition. We anticipate that the cross-scale code representations learned from transformer backbones could offer advantages over the fixed-scale features of mainstream convolutional codecs in these tasks.
Second, different automatic metrics for audio evaluation can produce inconsistent results, which is evidenced in our results. To further strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014). Despite this limitation, we provide a collection of demo speech samples publicly available in our codebase, which we hope will help demonstrate ESC’s performance and compensate for the absence of subjective metrics.
Besides, the primary focus of this work is to demonstrate the superiority of transformer and cross-scale frameworks over other mainstream methods, rather than to develop a production-ready codec like DAC or EnCodec. Nonetheless, given the scalability of transformers Kaplan et al. (2020), increasing the ESC model size and training it on larger and more diverse audio datasets also represent a promising direction for enhancing its practical applicability.
Finally, as discussed in Section 3.4, the cross-scale residual vector quantization (CS-RVQ) scheme requires the partial use of decoder layers during the encoding process, introducing additional latency as the bitrate increases. Similar to residual vector quantization, CS-RVQ requires careful sampling of the transmitted bitstream during training to achieve scalable bitrates within a single model. This sampling strategy can lead to performance trade-offs across different bitrates and may cause instability during training. Therefore, future research in speech foundational models could explore leveraging alternative recurrent structures Toderici et al. (2017); Johnston et al. (2018); Diao et al. (2020) to improve coding scalability and address these challenges.
@@ Appendix A Pre-training Paradigm
The proposed pre-training paradigm for optimizing vector quantization layers is detailed in Algorithm 3. During the pre-training phase, all vector quantization layers are bypassed, effectively reducing the codec to a standard autoencoder trained solely on reconstruction losses (Lines 2-4). Once the encoder and decoder reach a certain level of convergence, the VQ layers are reactivated, and joint optimization resumes. In the pre-training phase, we set the $\operatorname*{arg\,min}$ nearest-neighbor selection as an identity function, making $\bm{z}_{q}$ equal to the input vector $\bm{z}_{e}$.
@@
2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\mathcal{X}))$
3:     $\mathcal{L}=\mathcal{L}_{recon}(\mathcal{X},\hat{\mathcal{X}})$
@@ B.1 DAC Reproduction Setups
Our customized reproduction of DAC models closely follows the official development scripts. The original DAC model, designed for 16kHz audio signals, employs 12 VQ layers in its residual VQ module, supporting bitrates ranging from 0.5 kbps to 6.0 kbps. To ensure a fair comparison with ESC at similar bitrate levels, we extended the number of VQ layers in the RVQ module to 18, resulting in DAC-Base. For DAC-Tiny, we reduced the encoder dimension from 64 to 32 and the decoder dimension from 1536 to 288, while keeping other parameters unchanged. All DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our multilingual speech dataset. Additional configuration details can be found in the official release [^5] The official configuration for 16kHz DAC model is available at https://github.com/descriptinc/descript-audio-codec/blob/main/conf/final/16khz.yml.
@@ B.2 ESC Architecture Configurations
Overall, all three ESC variants are trained in distributed setups for 0.4 million iterations across 4 NVIDIA RTX 4090 GPUs with a total batch size of 36. These experiments took approximately 100 GPU hours.
B.2.1 Model Parameters
The parameter configurations for ESC-Base are provided in Table 3. For STFT transformation, we use a 20 ms window length and a 5 ms hop length, implemented with torchaudio. The number of FFT points is set to 382, resulting in a frequency dimension of 192. In the Swin Transformer, the layer depth represents the number of Swin Transformer Blocks (STBs) cascaded at each encoder and decoder layer. We use GELU activation functions and LayerNorm for normalization. In the down-sampling/up-sampling module, we use a scaling factor of $v=2$ to un-shuffle/shuffle along the frequency resolution only. Before the vector quantization layers, ESC processes two overlapping time frames together. To implement this, the flattened spectrum feature $\mathcal{Z}$ is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}$ to $\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$. Each frame is then split into sub-vectors, down-projected, and $l_{2}$ normalized before computing the distance matrix. The VQ layer at each bitstream of ESC-Base consumes $\log_{2}1024\times 3\times 150=4500$ bits per 3-second input speech (i.e., 1.5 kbps bitrate). For the scaled-up ESC-Large variant, we increase the STB layer depth from 2 to 4 while keeping the other configurations unchanged.
@@ B.2.1 Model Parameters
Table 3: Parameter configurations of model variant ESC-Base, which comprises 6 encoder/decoder layers.
B.2.2 Adversarial Training Setup
In the ESC-Base (adversarial) variant, the GAN discriminator is identical to the one used in DAC, consisting of a multi-period discriminator (MPD), multi-band discriminator (MBD), and multi-scale STFT discriminator (MSD), totaling over 42 million parameters. The adversarial loss formulation follows the official DAC-Base (adversarial) configuration. Additionally, we maintain the pre-training paradigm for 0.75 million iterations in this variant, with the discriminator intervening in training only after the pre-training stage finishes.
B.3 Details on Ablation Experiments
All ablation models operate on the complex STFT spectrum, as in ESC (SwinT + CS-RVQ), using the same STFT configurations specified in Table 3. These models were trained for 0.25 million iterations, with 0.025 million iterations allocated for pre-training. The Swin Transformer configurations mirror those used in ESC-Base. Similarly, the vector quantization setup in the CS-RVQ models follows that of ESC-Base. In total, the ablation experiments required approximately 80 hours on 4 RTX 4090 GPUs.
B.3.1 Convolution Blocks
For models with CNN backbones, the convolutional channel dimensions were set to match the hidden sizes of the STB-based models. Each CNN block consists of one residual unit and one downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency resolution only. The residual unit consists of two 2D convolutional layers, each followed by BatchNorm and Parametric ReLU activation.
B.3.2 Residual Vector Quantization Setups
For models using RVQs, we adapted the basic RVQ framework commonly used in time-domain codecs. To process frequency-domain spectrum features at the latent bottleneck, we combined RVQ with product vector quantization. Specifically, the flattened time frame vector is split into sub-group vectors, which are then recursively quantized, as in standard RVQs. We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5 kbps per bitstream, 6 in total).